OpenStudy (iwanttogotostanford):

A person goes to the store and buys 12 lemons, if they order 10 lemonades, and the lemonade is each 5.7 ounces, how many lemonades can be made?

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 } + \frac{ 3 }{ x } = \frac{ 1 }{ 2 }\] Unsolvable? Quite the contrary, George. \[\frac{ 3 }{ x } = \frac{ 1 }{ 2 } - \frac{ 4 }{ 5 }\] \[\frac{ 3 }{ x } = \frac{ 1(5) - 4(2) }{ 2(5) } = \frac{ 5 - 8 }{ 10 } = \frac{ -3 }{ 10 } \] \[\frac{ 3 }{ x } = \frac{ -3 }{ 10 }\] \[3\left( \frac{ 10 }{ -3 } \right) = x\] \[-10 = x\] As we can see, x = -10. George's conclusion was false because he had already assumed that x = 0.

OpenStudy (iwanttogotostanford):

oh wow, ok thank you! I get it now as my teacher did NOT explain very well.. sigh

OpenStudy (anonymous):

You're welcome ^_^ Note: I understand the sentiment. Sadly, you'll continue to have professors at higher education that don't explain too well.

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo I wil be having a couple more that I need explaining on if you're up for it!:-)

OpenStudy (anonymous):

Just post them on another thread, and we'll tackle it.

OpenStudy (iwanttogotostanford):

ok thank you

OpenStudy (anonymous):

Don't forget to close this one!